OpenStudy (anonymous):

Can anyone help me with this equation Salisbury High School decided to take their students on a field trip to a theme park. A total of 150 people went on the trip. Adults pay $45.00 for a ticket and students pay $28.50 for a ticket. How many students and how many adults went to the park if they paid a total of $4770?

OpenStudy (anonymous):

please write your full equation and how you got it

OpenStudy (anonymous):

Notice that there are two variables in this question, connected by two things. Number and money. So there should be two equations. x is number of adults, y the number of students. Number: x+y=150 Money: 45x+28.5y=4770 Do you require further help?

OpenStudy (anonymous):

yes how do you find the anwser

OpenStudy (anonymous):

Do you know how to solve simultaneous equations?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

confused

OpenStudy (anonymous):

help

OpenStudy (anonymous):

help

OpenStudy (anonymous):

is it : 160 adults and 44 kids

OpenStudy (anonymous):

is it???????????????

OpenStudy (anonymous):

sorry 106 adults and 44 kids

OpenStudy (anonymous):

lol sorry about that. Substituting back..nope. To solve this equation, I recommend rearranging the first eq into x=150-y then substituting into the second one, 45x+28.5y=4770 45(150-y)+28.5y=4770 that's how to solve simultaneous eqs.

OpenStudy (anonymous):

you should get more than one hundred students, about 120 of them.

OpenStudy (anonymous):

ohh now i got it: 125 students and 25 adults

OpenStudy (anonymous):

?

OpenStudy (anonymous):

A bit too much students though. Check your answers by substituting back into the equation.

OpenStudy (anonymous):

120 students and 30 adults

OpenStudy (anonymous):

????????

OpenStudy (anonymous):

yup! :) You've got it.

OpenStudy (anonymous):

i dont know how to solve the equation above 45(150-y) +28.5y=4770 After i reviewed my answers are different

OpenStudy (anonymous):

well, you do this, 45(150-y) +28.5y=4770 45*150 - 45y + 28.5y=4770 -16.5y=4770-45*150 y=120 Do you follow?

OpenStudy (anonymous):

OMG its too late. Multiplication off and then i go in and solve for x?